OpenStudy (anonymous):

find the angles between the pairs of vectors A = -2.00i + 6.00j and B = 2.00i - 3.00j

OpenStudy (anonymous):

@aum

OpenStudy (dumbcow):

\[\cos \theta = \frac{A*B}{|A| |B|}\] A*B is dot product

OpenStudy (anonymous):

the top gives me -22... ?

OpenStudy (anonymous):

right ?

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

so then what did I do wrong.. ? lol

OpenStudy (anonymous):

because thats not possible

OpenStudy (anonymous):

oh wait hold on im retarded

OpenStudy (dumbcow):

what about the bottom

OpenStudy (anonymous):

wait 180 isn't right... =/

OpenStudy (anonymous):

hello.. ?

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Why did you say something then leave... I hate those type, you are no help, BYE.